There are two common oxides of sulphur, one of which contains 50%O2 by weight, the other almost exactly 60%. The weights of sulphur which combine with 1 g of O2 are in the ratio of

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
                               FIRST OXIDE    SECOND OXIDE
Oxygen                 50%                      60%
sulphur(by diff)    50%                      40%
In first oxide 50 parts by weight of oxygen combine with sulphur=50 parts
1 part by weight of oxygen will combine with sulphur=50/50=1


In second oxide 60 parts by weight of oxygen combine with sulphur=40 parts
1 part by weight of oxygen will combine with sulphur=40/60=2/3


Compare the weights of sulphur which combine with same weight of oxygen in both oxides         
                                                           1:2/3
                                                           =3:2
 
Regards
Arun(askIITians forum expert)